# The Narrow Depth and Breadth of Corporate Responsible AI Research

### v1

@@ Abstract
The transformative potential of AI presents remarkable opportunities, but also significant risks, underscoring the importance of responsible AI development and deployment. Despite a growing emphasis on this area, there is limited understanding of industry’s engagement in responsible AI research, i.e., the critical examination of AI’s ethical, social, and legal dimensions. To address this gap, we analyzed over 6 million peer-reviewed articles and 32 million patent citations using multiple methods across five distinct datasets to quantify industry’s engagement. Our findings reveal that the majority of AI firms show limited or no engagement in this critical subfield of AI. We show a stark disparity between industry’s dominant presence in conventional AI research and its limited engagement in responsible AI. Leading AI firms exhibit significantly lower output in responsible AI research compared to their conventional AI research and the contributions of leading academic institutions. Our linguistic analysis documents a narrower scope of responsible AI research within industry, with a lack of diversity in key topics addressed. Our large-scale patent citation analysis uncovers a pronounced disconnect between responsible AI research and the commercialization of AI technologies, suggesting that industry patents rarely build upon insights generated by the responsible AI literature. This gap highlights the potential for AI development to diverge from a socially optimal path, risking unintended consequences due to insufficient consideration of ethical and societal implications. Our results highlight the urgent need for industry to publicly engage in responsible AI research to absorb academic knowledge, cultivate public trust, and proactively mitigate AI-induced societal harms.
@@ 1 Introduction
As AI continues to mediate our lives and relationships (Lazar,, 2022; Wagner et al.,, 2021), it brings a host of benefits (Taddeo and Floridi,, 2018), such as automation of routine tasks (Brynjolfsson et al.,, 2019), improved medical services (Coudray et al.,, 2018), and increased productivity (Noy and Zhang,, 2023). At the same time, many studies have demonstrated the ethical issues with AI deployment—including creating harm by recommending hateful content (Santos et al.,, 2021), generating discriminatory predictions (Buolamwini and Gebru,, 2018; Huszár et al.,, 2022; Dressel and Farid,, 2018), or sowing dissent through disinformation (Matz et al.,, 2017; Acquisti et al.,, 2015; Vlasceanu and Amodio,, 2022). Consequently, a general lack of trust in AI has emerged among the public (Mitre,, 2023). Hence, regulators, such as the US Congress, the White House, and the European Parliament, have joined scholars in questioning the actions of AI firms–firms that own one or more AI patents–and asking for more public engagement, including research (The White House, 2023b, ), on these societal issues (Kang,, 2023).
@@
In this article, we use the term “responsible AI” to denote the examination of the ethical and societal implications of AI. This includes improving its design, development and deployment, and assessing the appropriateness of AI use in specific situations (Barocas et al.,, 2020). Research into responsible AI is critical for aligning AI models with societal values (Van de Poel,, 2020), respecting fundamental human rights (Aizenberg and van den Hoven,, 2020), and contributing positively to society while minimizing potential harms (Hoffmann and Frase,, 2023; Smith,, 2023), and ensuring the trustworthiness of AI technologies (Knowles and Richards,, 2021). Therefore, firms creating these technologies need to critically and ethically evaluate their AI systems, as the benefits and costs of AI are not inherent to the technology but rather the result of how it is developed and deployed (Acemoglu,, 2021).
@@
AI firms are in a unique position in shaping the trajectory of AI development, possessing the resources, talent, and capability to drive innovation in this field (Ahmed et al.,, 2023; Frank et al.,, 2019). However, this privileged position also entails a heightened responsibility to ensure the socially beneficial and ethically sound development and deployment of AI (Rudner,, 1953; Johnson,, 2023; Bijker and Law,, 1994). Engaging in responsible AI research is crucial for AI firms to cultivate their absorptive capacity (Cohen and Levinthal,, 1990; Zahra and George,, 2002; Baruffaldi and Poege,, 2023), foster transparency and accountability (Brundage et al.,, 2020; Mittelstadt,, 2019), and navigate the complex trade-offs involved in aligning AI with societal values (Korinek and Balwit,, 2022). Public engagement and deliberation are essential in navigating the responsible development of AI, as the substantial conceptual ambiguities surrounding this field require input from diverse stakeholders (Raghavan,, 2023). Moreover, as AI firms are accountable for the consequences of their products, investing in responsible AI research is necessary to understand the limitations and potential risks associated with their technology (Holstein et al.,, 2019). Overall, prioritizing public engagement in responsible AI development not only benefits society as a whole but also provides tangible financial benefits to AI firms themselves by building trust among regulators and customers.
While research into responsible AI by firms could enable them to develop and deploy the technology in a socially beneficial manner (Askell et al.,, 2019), the extent and impact of industry involvement in responsible AI research remain ambiguous. The literature presents divergent perspectives on this topic, highlighting a tension between the potential benefits and risks of industry engagement in responsible AI. Critics argue that excessive industry involvement in AI research could compromise the integrity and independence of responsible AI initiatives (Abdalla and Abdalla,, 2021; Hao,, 2021; Young et al.,, 2022). For example, Baker and Hanna (2022) assert that this critical research stream is at risk, emphasizing the need for funding independent, grassroots-led research. This is in part because there is a significant concern that industry might have excessive influence over the trajectory of AI research (Young et al.,, 2022; Ahmed and Wahed,, 2020). Furthermore, some argue that industry collaborates with academics on responsible AI research to influence regulation, potentially undermining the objectivity of such efforts (Ochigame,, 2019). This perspective suggests that industry engagement in responsible AI research may be driven by self-interest rather than a genuine commitment to ethical AI development (Cath et al.,, 2018).
Conversely, new findings challenge the universally negative view of industry participation in responsible AI research. Emerging evidence indicates that certain firms have advanced significantly in this area (Zhang et al.,, 2022; Maslej et al.,, 2023), progressing past superficial contributions and implementation (de Laat,, 2021). This suggests that industry engagement in responsible AI research and practice is not always detrimental and that certain firms are committed to socially beneficial AI development. This tension in the literature underscores the need for further investigation into the nature and extent of industry’s engagement in responsible AI research.
@@
To answer these questions, we conducted a systematic analysis of more than 6 million peer-reviewed papers (2010-22) sourced from Scopus and 32 million USPTO patent citations (1947-2022) (Marx and Fuegi,, 2020). Our comprehensive dataset encompasses a wide range of conference and journal articles, adopting an expansive definition of responsible AI research. We employed multiple methods, including supervised machine learning and keyword search, to classify papers as responsible AI research. Leveraging five distinct datasets (see Table LABEL:table:3_distinct_data), we evaluated the extent of industry’s engagement in this important subfield. Our analyses consistently demonstrate that industry’s participation in responsible AI research lacks both depth and breadth.
@@
  • First, our comprehensive analysis of nearly all industry-authored AI papers reveals a concerning lack of industry engagement in responsible AI research. We employ two distinct methods—a rigorously tested supervised machine learning analysis and a keyword search approach—to classify responsible AI papers. The results suggest that nearly 90% of AI firms with commercial patents conduct no research into responsible AI. Even more alarming is the finding that among AI research firms—those with one or more AI publications—only 11.2% engage in substantive responsible AI research. Overall, our findings indicate a lack of meaningful industry engagement in responsible AI research.
  • Second, comparing industry to academic institutions by analyzing over 5.9 million papers, our study shows a significant disparity in the production of responsible AI research by leading AI firms. Leading AI firms produce significantly lower output in responsible AI research compared to their conventional AI research and the contributions of leading academic institutions.
  • Third, analyzing industry presence at AI conferences reveals a stark imbalance: while industry dominates conventional AI conferences, its engagement in responsible AI conferences is limited, highlighting a lack of commitment to responsible AI advancement.
  • Fourth, by employing various natural language processing (NLP) methods on paper abstracts, we find that AI firms tend to have a more limited research focus on responsible AI compared to academic researchers. Specifically, areas such as moral considerations, environmental concerns, and ethical consequences receive less attention in industry papers.
  • Finally, to the best of our knowledge, this work is the first empirical evidence that shows that responsible AI research has a limited impact on industry’s commercial inventions. Our comprehensive analysis of over 32 million patent citations reveals a notable gap between responsible AI research and its application by industry, suggesting limited commercial integration of responsible AI research.
Our findings indicate that AI’s rapid commercialization is proceeding with insufficient attention to its potential consequences, despite mounting pressure from regulators (The White House, 2023a, ), activists (Krafft et al.,, 2021), and employees (Ahmed and Jia,, 2023; Belfield,, 2020) to prioritize responsible development. Industry’s limited engagement in responsible AI research is concerning, given its leading role in AI development. Our analysis shows that AI firms have minimal involvement in responsible AI research and development, as evidenced by their publications and patent citations. This suggests that the current trajectory of AI may not be socially optimal.
@@ 2 Leading
data (Marx and Fuegi,, 2020)
@@ 2 Related Literature on Industry’s Engagement in Responsible AI Research
Recent research has documented industry’s increasing dominance of the frontiers of AI research (Frank et al.,, 2019; Koch et al.,, 2021; Ahmed et al.,, 2023). Studies have shown that industry not only has an increased presence but also influences the trajectory of AI research by creating the majority of large AI models (Benkler,, 2019; Jurowetzki et al.,, 2021; Ahmed et al.,, 2023). Concerns arise from the significant influence industry may exert over the direction of AI research (Young et al.,, 2022; Ahmed and Wahed,, 2020), including hiring academics (Ahmed et al.,, 2023; Simonite,, 2020; Gofman and Jin,, 2023) and directly funding AI research at universities (Abdalla and Abdalla,, 2021; Hao,, 2021; Young et al.,, 2022). Consequently, scholars worry that academia is falling behind (Ahmed and Wahed,, 2020; Jurowetzki et al.,, 2021), with industry taking the lead in pushing the boundaries of AI research and development.
However, there is conflicting evidence regarding industry’s engagement in responsible AI research. One stream of literature suggests that engagement in responsible AI research has increased notably (Maslej et al.,, 2023, 2024). The Stanford AI Index 2022 reports a 71% year-over-year increase in industry co-authored publications at leading responsible AI conferences in recent years (Zhang et al.,, 2022). Research has also suggested that certain AI firms have genuinely engaged with their responsible AI practices (de Laat,, 2021). In fact, concerns have grown to such an extent that scholars have predicted AI firms might shape the future research agenda in the field of responsible AI (Hao,, 2021; Young et al.,, 2022; Baker and Hanna,, 2022; Jazwinska,, 2022). This has led some to argue that industry collaborates with academics on responsible AI research to sway regulation, potentially compromising the objectivity of these efforts (Ochigame,, 2019; Cath et al.,, 2018). Moreover, critics contend that excessive industry involvement in AI research could compromise the integrity and independence of responsible AI initiatives (Abdalla and Abdalla,, 2021; Hao,, 2021; Young et al.,, 2022; Baker and Hanna,, 2022).
On the other hand, another stream of literature predicts and argues that industry might have limited incentive to engage in the development of responsible AI (Askell et al.,, 2019). The “ethics washing” literature claims that firms create the illusion of ethical behavior without substantive work in responsible AI (de Laat,, 2021; Rakova et al.,, 2021; Falco et al.,, 2021). This suggests that industry has limited engagement in both responsible AI research and practice, contradicting prior claims.
Scholars have also examined the breadth of industry’s responsible AI research, focusing on the range of topics addressed. However, these examinations have been limited in scale. Recent textual analyses indicate that while certain areas of responsible AI have gained traction over time (Laufer et al.,, 2022), mainstream AI research often prioritizes the interests of large corporations over societal needs, potentially overlooking harmful consequences (Birhane et al.,, 2022). It is important to note that a significant drawback of these studies is their narrow focus on only the most prestigious responsible AI conferences, which may not provide a comprehensive view of the field.

### v2

@@ Abstract
The transformative potential of AI presents remarkable opportunities, but also significant risks, underscoring the importance of responsible AI development and deployment. Despite a growing emphasis on this area, there is limited understanding of industry’s engagement in responsible AI research, i.e., the systematic examination of AI’s ethical, social, and legal dimensions. To address this gap, we analyzed over 6 million peer-reviewed articles and 32 million patent citations using multiple methods across five distinct datasets to quantify industry’s engagement. Our analysis reveals notable heterogeneity between industry’s substantial presence in conventional AI research and its comparatively modest engagement in responsible AI. Leading AI firms exhibit significantly lower output in responsible AI research compared to their conventional AI research and the contributions of leading academic institutions. Our linguistic analysis reveals a more concentrated scope of responsible AI research within industry, with fewer distinct key topics addressed. Our large-scale patent citation analysis uncovers limited linkage between responsible AI research and the commercialization of AI technologies, suggesting that industry patents infrequently draw upon insights from the responsible AI literature. These patterns raise important questions about the integration of responsible AI considerations into commercialization practices, with potential implications for the alignment of AI development with broader societal objectives. Our results highlight the need for industry to publicly engage in responsible AI research to absorb academic knowledge, cultivate public trust, and proactively address the societal dimensions of AI development.
@@ 1 Introduction
As AI continues to mediate our lives and relationships (Lazar, 2022; Wagner, Strohmaier, Olteanu, Kıcıman, Contractor, and Eliassi-Rad, 2021), it brings a host of benefits (Taddeo and Floridi, 2018), such as automation of routine tasks (Brynjolfsson, Hui, and Liu, 2019), improved medical services (Coudray, Ocampo, Sakellaropoulos, Narula, Snuderl, Fenyö, Moreira, Razavian, and Tsirigos, 2018), and increased productivity (Noy and Zhang, 2023). At the same time, studies have documented ethical challenges associated with AI deployment, including algorithmic recommendations of harmful content (Santos, Lelkes, and Levin, 2021), discriminatory predictions (Buolamwini and Gebru, 2018; Huszár, Ktena, O’Brien, Belli, Schlaikjer, and Hardt, 2022; Dressel and Farid, 2018), and the spread of disinformation (Matz, Kosinski, Nave, and Stillwell, 2017; Acquisti, Brandimarte, and Loewenstein, 2015; Vlasceanu and Amodio, 2022). Consequently, public discourse around AI has increasingly focused on questions of accountability and oversight (Mitre, 2023; Haenlein, Huang, and Kaplan, 2022). Regulators, including the US Congress, the White House, and the European Parliament, have called for greater transparency and research engagement from AI firms (organizations that own one or more AI patents) on these societal dimensions (The White House, 2023b; Kang, 2023).
@@
In this article, we use the term “responsible AI” to denote the examination of the ethical and societal implications of AI. This includes improving its design, development and deployment, and assessing the appropriateness of AI use in specific contexts (Barocas, Biega, Fish, Niklas, and Stark, 2020). Research into responsible AI contributes to aligning AI models with societal values (Van de Poel, 2020), respecting fundamental human rights (Aizenberg and van den Hoven, 2020), and understanding how to maximize societal benefits while addressing potential harms (Hoffmann and Frase, 2023; Smith, 2023), as well as establishing the trustworthiness of AI technologies (Knowles and Richards, 2021). Firms creating these technologies are positioned to critically evaluate their AI systems, as the outcomes of AI depend substantially on how it is developed and deployed (Acemoglu, 2021).
@@
AI firms are in a unique position in shaping the trajectory of AI development, possessing the resources, talent, and capability to drive innovation in this field (Ahmed, Wahed, and Thompson, 2023; Frank, Wang, Cebrian, and Rahwan, 2019). This position affords opportunities for these organizations to contribute to the socially beneficial development and deployment of AI (Rudner, 1953; Johnson, 2023; Bijker and Law, 1994). Engagement in responsible AI research may enable firms to develop absorptive capacity (Cohen and Levinthal, 1990; Zahra and George, 2002; Baruffaldi and Poege, 2023), foster transparency and accountability (Brundage, Avin, Wang, Belfield, Krueger, Hadfield, Khlaaf, Yang, Toner, Fong, Maharaj, Koh, Hooker, Leung, Trask, Bluemke, Lebensold, O’Keefe, Koren, and Ryffel, 2020; Mittelstadt, 2019), and navigate the trade-offs inherent in aligning AI systems with societal values (Korinek and Balwit, 2022). Public engagement and deliberation are essential in navigating the responsible development of AI, as the substantial conceptual ambiguities surrounding this field require input from diverse stakeholders (Raghavan, 2023). Given that AI firms bear accountability for the societal consequences of their products, engagement in responsible AI research offers a pathway toward understanding the limitations and risks associated with their technologies (Holstein, Wortman Vaughan, Daumé, Dudik, and Wallach, 2019). Overall, prioritizing public engagement in responsible AI development not only benefits society as a whole but also provides tangible financial benefits to AI firms themselves by building trust among regulators and customers.
While research into responsible AI by firms could enable them to develop and deploy the technology in a socially beneficial manner (Askell, Brundage, and Hadfield, 2019), the extent and nature of industry involvement in responsible AI research remain underexplored. The literature presents divergent perspectives on this topic. One strand of scholarship examines the relationship between industry participation and the independence of responsible AI research (Abdalla and Abdalla, 2021; Hao, 2021; Young, Katell, and Krafft, 2022). For example, Baker and Hanna (2022) highlight the role of funding for independent, grassroots-led research. Related work analyzes how industry’s substantial presence in AI research may shape the trajectory of the field (Young, Katell, and Krafft, 2022; Ahmed and Wahed, 2020). Other scholars have examined industry collaborations with academics on responsible AI research in relation to regulatory dynamics (Ochigame, 2019), exploring the various motivations that may underlie such engagement (Cath, Wachter, Mittelstadt, Taddeo, and Floridi, 2018).
At the same time, emerging findings present a more nuanced view of industry participation in responsible AI research. Recent evidence indicates that certain firms have made notable progress in this area (Zhang, Maslej, Brynjolfsson, Etchemendy, Lyons, Manyika, Ngo, Niebles, Sellitto, Sakhaee, Yoav Shoham, and Perrault, 2022; Maslej, Fattorini, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Ngo, Niebles, Parli, Shoham, Wald, Clark, and Perrault, 2023; Ahmed and Jia, 2023), advancing beyond initial efforts (de Laat, 2021). These findings suggest that industry engagement in responsible AI research exhibits considerable heterogeneity, with some organizations engaged in socially beneficial AI development. This heterogeneity motivates systematic empirical investigation into the nature and extent of industry’s engagement in responsible AI research.
@@
To answer these questions, we conducted a systematic analysis of more than 6 million peer-reviewed papers (2010-22) sourced from Scopus and 32 million USPTO patent citations (1947-2022) (Marx and Fuegi, 2020). Our comprehensive dataset encompasses a wide range of conference and journal articles, adopting an expansive definition of responsible AI research. We employed multiple methods, including supervised machine learning and keyword search, to classify papers as responsible AI research. Leveraging five distinct datasets (see Table LABEL:table:3_distinct_data), we evaluated the extent of industry’s engagement in this important subfield. Our analyses reveal considerable heterogeneity in industry’s engagement with responsible AI research, with most firms exhibiting limited depth and breadth of participation relative to their investments in conventional AI development.
@@
  • First, our comprehensive analysis of industry-authored AI papers documents patterns of industry engagement in responsible AI research. We employ multiple complementary methods, including a supervised machine learning classifier, a keyword-based approach, and large language model classification (Gemini), to identify responsible AI papers. Among AI research firms (those with one or more AI publications), we observe that 34.5% engage in responsible AI research. These findings provide an empirical baseline for understanding the current landscape of industry participation in this emerging subfield.
  • Second, comparing industry to academic institutions by analyzing over 5.9 million papers, our study documents differences in the production of responsible AI research. Leading AI firms exhibit lower output in responsible AI research relative to their conventional AI research and relative to the contributions of leading academic institutions. At the same time, we observe that some leading firms have published responsible AI research at levels comparable to elite academic institutions, suggesting heterogeneity across firms.
  • Third, analyzing industry presence at AI conferences reveals differential patterns: while industry has substantial presence at conventional AI conferences, its participation at responsible AI conferences is comparatively limited.
  • Fourth, by employing various natural language processing (NLP) methods on paper abstracts, we find that AI firms’ responsible AI research tends to focus on different topics compared to academic researchers. Areas such as moral considerations, environmental concerns, and ethical consequences appear less frequently in industry papers.
  • Finally, to the best of our knowledge, this work provides the first large-scale empirical documentation of the relationship between responsible AI research and industry’s commercial inventions. Our analysis of over 32 million patent citations reveals that industry patents infrequently cite responsible AI research, suggesting limited integration of this research stream into commercial development.
Our findings indicate that AI’s rapid commercialization is proceeding with modest attention to its potential consequences, despite mounting pressure from regulators (The White House, 2023a), civil society (Krafft, Young, Katell, Lee, Narayan, Epstein, Dailey, Herman, Tam, Guetler, Bintz, Raz, Jobe, Putz, Robick, and Barghouti, 2021), and employees (Ahmed and Jia, 2023; Belfield, 2020) for greater attention to responsible development. Given industry’s leading role in AI development, understanding the extent of corporate engagement in responsible AI research has important implications for both scholarship and policy. Our analysis indicates that AI firms have comparatively limited involvement in responsible AI research and development, as measured by publications and patent citations. These patterns raise questions about whether current market incentives are sufficient to align corporate AI development with broader societal considerations.
@@ 2 Leading
data (Marx and Fuegi, 2020)
@@ 2 Related Literature on Industry’s Engagement in Responsible AI Research
Recent research has documented industry’s growing presence at the frontiers of AI research (Frank, Wang, Cebrian, and Rahwan, 2019; Koch, Denton, Hanna, and Foster, 2021; Ahmed, Wahed, and Thompson, 2023). Studies indicate that industry not only maintains an increased presence but also shapes the trajectory of AI research through developing the majority of large AI models (Benkler, 2019; Jurowetzki, Hain, Mateos-Garcia, and Stathoulopoulos, 2021; Ahmed, Wahed, and Thompson, 2023). Scholars have examined the implications of industry’s influence over the direction of AI research (Young, Katell, and Krafft, 2022; Ahmed and Wahed, 2020), including patterns of hiring academics (Ahmed, Wahed, and Thompson, 2023; Simonite, 2020; Gofman and Jin, 2023) and funding AI research at universities (Abdalla and Abdalla, 2021; Hao, 2021; Young, Katell, and Krafft, 2022). This body of work has raised questions about the evolving division of labor between academia and industry in AI research (Ahmed and Wahed, 2020; Jurowetzki, Hain, Mateos-Garcia, and Stathoulopoulos, 2021).
However, the literature presents mixed findings regarding industry’s engagement in responsible AI research. One stream of research suggests that such engagement has increased notably (Maslej, Fattorini, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Ngo, Niebles, Parli, Shoham, Wald, Clark, and Perrault, 2023, 2024). The Stanford AI Index 2022 reports a 71% year-over-year increase in industry co-authored publications at leading responsible AI conferences (Zhang, Maslej, Brynjolfsson, Etchemendy, Lyons, Manyika, Ngo, Niebles, Sellitto, Sakhaee, Yoav Shoham, and Perrault, 2022). Research has also documented cases where AI firms have substantively engaged with responsible AI practices (de Laat, 2021). Building on these observations, some scholars have examined whether AI firms might shape the future research agenda in responsible AI (Hao, 2021; Young, Katell, and Krafft, 2022; Baker and Hanna, 2022; Jazwinska, 2022). Related work has explored the relationship between industry-academic collaboration in responsible AI research and regulatory processes (Ochigame, 2019; Cath, Wachter, Mittelstadt, Taddeo, and Floridi, 2018). Patent citation analysis suggests that firms’ responsible AI publications may serve regulatory or reputational functions rather than directly informing product development (Ahmed, Polidoro Jr, and Das, 2026). Some scholars have also examined how varying levels of industry involvement may interact with the independence of responsible AI research initiatives (Abdalla and Abdalla, 2021; Hao, 2021; Young, Katell, and Krafft, 2022; Baker and Hanna, 2022).
A separate stream of literature examines whether industry may have limited incentives to engage substantively in responsible AI development (Askell, Brundage, and Hadfield, 2019). Some scholars have characterized certain forms of industry engagement as primarily symbolic rather than substantive, a phenomenon sometimes termed "ethics washing" (de Laat, 2021; Rakova, Yang, Cramer, and Chowdhury, 2021; Falco, Shneiderman, Badger, Carrier, Dahbura, Danks, Eling, Goodloe, Gupta, Hart, Jirotka, Johnson, Lapointe, Llorens, Mackworth, Maple, Pálsson, Pasquale, Alan, and Yeong, 2021). This perspective suggests that observed industry engagement in responsible AI may vary in depth and substance, presenting a different interpretation than the literature documenting increased engagement.
Scholars have also examined the topical breadth of industry’s responsible AI research. Prior studies, while offering valuable insights, have typically focused on specific venues or time periods. Recent textual analyses indicate that certain areas of responsible AI have gained traction over time (Laufer, Jain, Cooper, Kleinberg, and Heidari, 2022), and that the thematic priorities of mainstream AI research may differ from topics emphasized in academic responsible AI scholarship (Birhane, Kalluri, Card, Agnew, Dotan, and Bao, 2022). One limitation of prior studies is their focus on the most prestigious responsible AI conferences, which may not fully capture the breadth of research activity in the field. Our study addresses this gap by employing a more comprehensive sampling strategy across multiple data sources.
Industry engagement in responsible AI research exhibits notable heterogeneity, with certain firms producing scholarship at levels comparable to leading academic institutions. For example, Anthropic, founded in 2021 with safety as a central organizational mission, has published work on "Constitutional AI," a method for training AI systems using explicit principles rather than solely human feedback (Bai, Kadavath, Kundu, Askell, Kernion, Jones, Chen, Goldie, Mirhoseini, McKinnon, and others, 2022). Anthropic’s other works have contributed to interpretability research examining neural network representations through dictionary learning methods (Templeton, Conerly, Marcus, Lindsey, Bricken, Chen, Pearce, Citro, Ameisen, Jones, Cunningham, Turner, McDougall, MacDiarmid, Tamkin, Durmus, Hume, Mosconi, Freeman, Sumers, Rees, Batson, Jermyn, Carter, Olah, and Henighan, 2024). Similarly, Google DeepMind maintains one of the larger industry research portfolios in this area, with published work on frontier safety frameworks for managing risks from advanced AI systems, holistic safety evaluation methodologies, and interpretability tools such as Gemma Scope (Lieberum, Rajamanoharan, Conmy, Smith, Sonnerat, Varma, Kramár, Dragan, Shah, and Nanda, 2024). These examples indicate that substantive responsible AI research from industry does occur. However, less is known about the prevalence and characteristics of such engagement across the broader population of AI firms, a gap our empirical analysis seeks to address.